A cylinder with rotational inertia

I1 = 3.2 kg · m2

rotates clockwise about a vertical axis through its center with angular speed

ω1 = 5.8 rad/s.

A second cylinder with rotational inertia

I2 = 1.2 kg · m2

rotates counterclockwise about the same axis with angular speed

ω2 = 6.2 rad/s.

If the cylinders couple so they have the same rotational axis, what is the angular speed of the combination (in rad/s)?

What percentage of the original kinetic energy is lost to friction?
Answer #1

solution ₃ I Feven the confervation af angular momentum (I + I 2) 2) = I, J, + IzQ2 3 d = 1, D1 + 12 Q2 I + I2 = 3.94.500 + 1